OpenStudy (anonymous):

What is the derivative of 2-(1/x) and x-3?

OpenStudy (anonymous):

D[2-(1/x)] = D[2] - D[1/x] = 0 - D[x^-1] = -(-x^-2) = x^2 You should be able to do the other one. If you need more help, let me know.